OpenStudy (celinegirl):

hi please check my answer (1/5) The graph of f(x) = (0.5)^x is replaced by the graph of g(x) = (0.5)^x − k. If g(x) is obtained by shifting f(x) down by 2 units, the value of k is ________. k = 7

OpenStudy (trojanpoem):

Why 7 ?

OpenStudy (trojanpoem):

let f(x) be a function to shift it up "a" units f(x) = f(x) + a to shift it down "a" units: f(x) = f(x) - a

OpenStudy (celinegirl):

bc a friend helped me get the answer i just don't know if its 100% correct.. it could be -7

OpenStudy (trojanpoem):

k = 2

OpenStudy (celinegirl):

hm idk

OpenStudy (trojanpoem):

https://www.desmos.com/calculator f\left(x\right)\ =\ \left(0.5\right)^x g\left(x\right)\ =\ \left(0.5\right)^x-2

OpenStudy (trojanpoem):

Shift it down means, subtracting from the y

OpenStudy (trojanpoem):

say you have y = x

OpenStudy (celinegirl):

idk i still think its 7 sorry